A beauty product company is developing a new fragrance named Happy Forever. There is a probability of 0.47 that consumers will love Happy Forever, and in this case, annual sales will be 1.08 million bottles; a probability of 0.39 that consumers will find the smell acceptable and annual sales will be 220,000 bottles; and a probability of 0.14 that consumers will find the smell unpleasant and annual sales will be only 50,000 bottles. The selling price is $39, and the variable cost is $10 per bottle. Fixed production costs will be $1.01 million per year, and depreciation will be $1.18 million. Assume that the marginal tax rate is 40 percent. What are the expected annual incremental after-tax free cash flows from the new fragrance?

Annual Incremental Cash Flows $___________
Answer #1

rate positively ... let me know if you need any clarification..

love Happy Forever smell acceptable smell unpleasant
i Sales                  1,080,000                 220,000                   50,000
ii Selling price                               39                          39                           39
iii=i*ii Total sales                42,120,000             8,580,000              1,950,000
iv Variable cost                10,800,000             2,200,000                 500,000
v Fixed cost                  1,010,000             1,010,000              1,010,000
vi Depreciation                  1,180,000             1,180,000              1,180,000
vii=iv+v+vi Total cost                12,990,000             4,390,000              2,690,000
vii=iii-iv Profit before tax                29,130,000             4,190,000               (740,000)
viii=vii*40% Tax @ 40%                11,652,000             1,676,000               (296,000)
ix=vii-viii Profit after tax                17,478,000             2,514,000               (444,000)
x=ix+vi Cash flow                18,658,000             3,694,000                 736,000
xi Probability =                            0.47                       0.39                       0.14
xii=xi*x expected annual cash flow                  8,769,260             1,440,660                 103,040 10,312,960
Ans =                10,312,960